Could someone please show me how to multiply and then simplify this expression. I have a few other problems like it and I am not sure how to solve it.
sqrt 13 * sqrt 5
I understand that by multiping 13 * 5= sqrt 65
How do I simplify this????

Answer by stanbon(75887)

You cannot simplify it because 65 has no perfect square factors.
--------------
It would be different with sqrt(3)*sqrt(6)= sqrt(18)
18 = 9*2, so you can simplify sqrt(18)
-------
sqrt(18) = sqrt(9*2) = 3sqrt2
===============
Cheers,
Stan H.